The slope of a line is -16 and the point (10, 18) is on the line. Express the line in slope-intercept form.
Answer by checkley77(12844)

The slope of a line is -16 and the point (10, 18)
Y=mX+b
18=-16*10+b
18=-160+b
b=18+160
b=178 the Y intercept.
Y=-16X+178 ans.